# PoseMamba: Monocular 3D Human Pose Estimation with Bidirectional Global-Local Spatio-Temporal State Space Model

### v1

@@ Introduction
  • To the best of our knowledge, we are the first to propose the pure Mamba-based model, PoseMamba, for 3D HPE under the category of 2D-to-3D lifting.
@@
Existing 3D human pose estimation methods can be categorized through two perspectives. Firstly, these methods can be divided into two types based on the input video type: multi-view and monocular approaches. Approaches that depend on multi-view inputs (Zhang et al. 2021; Reddy et al. 2021; Chun, Park, and Chang 2023) require multiple cameras capturing different perspectives, which may pose challenges in practical applications. Secondly, these methods can be divided into direct 3D HPE methods and 2D-3D lifting methods. Direct 3D HPE methods (Pavlakos, Zhou, and Daniilidis 2018; Sun et al. 2018; Zhou et al. 2019) derive the spatial coordinates of joints directly from video frames without intermediary steps. In contrast, 2D-3D lifting methods first employ readily available 2D pose detectors (Chen et al. 2018; Sun et al. 2019; Newell, Yang, and Deng 2016) before elevating 2D coordinates to 3D space (Zhao et al. 2023; Zhu et al. 2023; Zhang et al. 2022). However, existing 3D human pose estimation methods mostly use CNN or Transformer as the fundamental model. These CNN-based and Transformer-based architectures have inherent shortcomings: CNN has a limited receptive field which hinders their ability to capture broader spatial information, and Transformers demand high computational resources, resulting in expensive training and deployment. In this paper, we adopt SSM for 2D-3D lifting to learn global-local spatial-temporal information with linear complexity.
@@
Recently, Mamba (Gu and Dao 2023) has achieved a significant breakthrough with its linear-time inference and efficient training methodology. Building on the success of Mamba, MoE-Mamba (Pióro et al. 2024) amalgamated Mixture of Experts with Mamba, unlocking the scalability potential of SSMs and achieving performance akin to Transformers. For vision applications, Vision Mamba (Liu et al. 2024) and VMamba (Zhu et al. 2024) employed bidirectional SSM blocks and the cross-scan module, respectively, to enhance data-dependent global visual context. However, the exploration of Mamba’s potential in 3D human pose estimation remains untapped. In this paper, we do not simply apply SSM to pose estimation. We compare unidirectional scanning with bidirectional scanning and observe inaccuracies in limb recognition. Unlike Vision Mamba and VMamba, we enhance the spatial scanning method for 3D human pose estimation and propose bidirectional global-local spatial-temporal scanning to learn global-local spatial-temporal correlation sufficiently.
@@
$\begin{split}&\dot{h}(t)=\bm{A}h(t)+\bm{B}u(t)\\ &y(t)=\bm{C}h(t)+{D}x(t)\end{split}$ (1)
@@
Experiment
Datasets and Evaluation Metrics
Implementation Details
Model Variants
Method $N$ $d_{m}$ $T$ Params MACs
PoseMamba-S 20 64 243 0.860 M 3.587 G
PoseMamba-B 20 128 243 3.358 M 13.943 G
PoseMamba-L 40 128 243 6.714 M 27.881 G
Experimental settings
Performance comparison on Human3.6M
@@
Performance comparison on MPI-INF-3DHP

### v2

@@ Introduction
  • To the best of our knowledge, we are the first to introduce a novel bidirectional global-local spatio-temporal modeling approach and logical geometric scanning strategy within the Mamba framework, PoseMamba, for 3D HPE under the category of 2D-to-3D lifting.
@@
Existing 3D human pose estimation methods can be categorized through two perspectives. Firstly, these methods can be divided into two types based on the input video type: multi-view and monocular approaches. Approaches that depend on multi-view inputs (Zhang et al. 2021; Reddy et al. 2021; Chun, Park, and Chang 2023) require multiple cameras capturing different perspectives, which may pose challenges in practical applications. Secondly, these methods can be divided into direct 3D HPE methods and 2D-3D lifting methods. Direct 3D HPE methods (Pavlakos, Zhou, and Daniilidis 2018; Sun et al. 2018; Zhou et al. 2019; Huang et al. 2023) derive the spatial coordinates of joints directly from video frames without intermediary steps. In contrast, 2D-3D lifting methods first employ readily available 2D pose detectors (Chen et al. 2018; Sun et al. 2019; Newell, Yang, and Deng 2016) before elevating 2D coordinates to 3D space (Zhao et al. 2023; Zhu et al. 2023; Zhang et al. 2022). Existing works (Holmquist and Wandt 2023; Shan et al. 2023) use multi-hypothesis approach to improve depth ambiguity in 3D HPE. DSED (Liu et al. 2022) addresses the self-occlusion problem in 3D HPE by explicitly reasoning about occlusion relationships in multi-person scenarios. While HumMUSS (Mondal, Alletto, and Tome 2024) first explores bidirectional SSM modeling for human motion understanding, our work introduces a novel bidirectional global-local spatio-temporal approach and logical geometric scanning strategy tailored for 3D HPE. PoseMagic (Zhang et al. 2024) introduces a hybrid Mamba-GCN architecture, but its reliance on GCN for capturing local details may lead to insufficient detail for complex actions. In contrast, our PoseMamba captures local movement details more effectively and improves performance through its bidirectional global-local spatio-temporal modeling method and logical geometric scanning strategy.
@@
Recently, Mamba (Gu and Dao 2023) has achieved a significant breakthrough with its linear-time inference and efficient training methodology. Building on the success of Mamba, MoE-Mamba (Pióro et al. 2024) amalgamated Mixture of Experts with Mamba, unlocking the scalability potential of SSMs and achieving performance akin to Transformers. For vision applications, Vision Mamba (Zhu et al. 2024) and VMamba (Liu et al. 2024) employed bidirectional SSM blocks and the cross-scan module, respectively, to enhance data-dependent global visual context. However, the exploration of Mamba’s potential in 3D human pose estimation remains untapped. In this paper, we do not simply apply SSM to pose estimation. We compare unidirectional scanning with bidirectional scanning and observe inaccuracies in limb recognition. Unlike Vision Mamba and VMamba, we enhance the spatial scanning method for 3D human pose estimation and propose bidirectional global-local spatial-temporal scanning to learn global-local spatial-temporal correlation sufficiently.
@@
$\begin{split}&\dot{h}(t)=\bm{A}h(t)+\bm{B}x(t)\\ &y(t)=\bm{C}h(t)+{D}x(t)\end{split}$ (1)
@@
PoseMamba-X 243 $\times$ 26.5 M 109.9 G 452 M 37.1/31.5 14.8
Experiment
Datasets and Evaluation Metrics
Implementation Details
Model Variants
Method $N$ $d_{m}$ $T$ Params MACs
PoseMamba-S 20 64 243 0.860 M 3.587 G
PoseMamba-B 20 128 243 3.358 M 13.943 G
PoseMamba-L 40 128 243 6.714 M 27.881 G
PoseMamba-X 40 256 243 26.535 M 109.909 G
@@
Experimental settings
Performance comparison on Human3.6M
Performance comparison on MPI-INF-3DHP
@@ Conclusion
Acknowledgements. This work was supported by the National Natural Science Foundation of China under Grant 62406120 and the Guangxi Science and Technology Project (GuiKe-AB21196034).